# Extensions of the Busemann-Petty Problem for Arbitrary Measures

### v1

@@ 1. Introduction
This interesting problem triggered a series of deep developments connecting convex geometry and Fourier analysis. The answer was shown to be affirmative for dimensions $n\leq 4$ and negative for $n\geq 5$, after major contributions culminating in the analytic characterization of intersection bodies due to Koldobsky (see, e.g., [BusemannPetty1956, Gardner1994, Koldobsky98b, Koldobsky98]).
@@
The Busemann-Petty problem for arbitrary measures was addressed by Zvavitch in [zvavitch2005busemann], where two different measures were considered, one for measuring the bodies, and a separate one for the sections. We state here the theorem by Zvavitch (see Section 2 below for the definition of positive-definiteness).
@@
This paper extends Zvavitch’s approach in two directions. First, we observe that a very mild modification in the argument allows the monotonicity assumption of the function in (2) to be reversed (so that a function previously required to be non-increasing may instead be taken non-decreasing), and, more generally, to be decomposed as a sum of a radially non-decreasing part and a radially non-increasing part. Second, through a modification of hypothesis (3), we permit different measurement schemes for each body: each may be measured with its own pair of densities (one for (n1)(n-1)-dimensional sections and one for $n$-dimensional volume). More precisely, $K$ is measured with the densities $f_{n-1}$ and $f_{n}$, while $L$ is measured with $g_{n-1}$ and $g_{n}$. It is worth pointing out that, in this direction, an isomorphic version of the Busemann–Petty problem with one measure assigned to each body was considered in [koldobsky2019measure, Theorem 1], where a comparison between measures was established in terms of the outer volume ratio distance to the class of generalized intersection bodies.
@@ 2. Preliminaries
We follow [GelfandVilenkin, KoldobskyBook] for standard definitions concerning the Fourier transform of distributions. We denote by $\mathcal{S}$ the Schwartz space of complex-valued test functions on $\mathbb{R}^{n}$—that is, $C^{\infty}$ functions decreasing faster than any polynomial together with all their derivatives.
@@
A distribution $f$ is called positive if $\langle f,\varphi\rangle\geq 0$ for every nonnegative $\varphi\in\mathcal{S}$. It is called positive-definite if $\hat{f}$ is positive. Moreover, any positive-definite distribution is the Fourier transform of a nonnegative tempered measure (see [GelfandVilenkin, KoldobskyBook]).
@@
The Fourier transform of an even homogeneous distribution of degree $p\in(-n,0)$ is again an even homogeneous distribution of degree np-n-p (see [KoldobskyBook]).
Let $f$ be a homogeneous function of degree $n+p\in(-n,0)$ which is also a positive-definite distribution. Then the Fourier transform $\hat{f}$ is a tempered measure that can be expressed in polar coordinates (see [KoldobskyBook, Lemma 2.24]). Hence there exists a nonnegative measure $\gamma_{0}$ on $S^{n-1}$ such that
@@
The following proposition, known as the spherical Parseval formula, corresponds to [KoldobskyBook, Corollary 3.23] with $k=1$.
@@
The relation between the Fourier and Radon transforms is given by (see [KoldobskyBook, Lemma 3.7])

### v2

@@ 1. Introduction
This interesting problem triggered a series of deep developments connecting convex geometry and Fourier analysis. The answer was shown to be affirmative for dimensions $n\leq 4$ and negative for $n\geq 5$, after major contributions culminating in the analytic characterization of intersection bodies. See [1, 2, 4, 5, 6] for the solution of the Busemann-Petty problem, and its history.
@@
The Busemann-Petty problem for arbitrary measures was addressed by Zvavitch in [8], where two different measures were considered, one for measuring the bodies, and a separate one for the sections. We state here the theorem by Zvavitch (see Section 2 below for the definition of positive-definiteness).
@@
This paper extends Zvavitch’s approach in two directions. First, we observe that a very mild modification in the argument allows the monotonicity assumption of the function in (2) to be reversed (so that a function previously required to be non-increasing may instead be taken non-decreasing), and, more generally, to be decomposed as a sum of a radially non-decreasing part and a radially non-increasing part. Second, through a modification of hypothesis (3), we permit different measurement schemes for each body: each may be measured with its own pair of densities (one for (n1)(n-1)-dimensional sections and one for $n$-dimensional volume). More precisely, $K$ is measured with the densities $f_{n-1}$ and $f_{n}$, while $L$ is measured with $g_{n-1}$ and $g_{n}$. It is worth pointing out that, in this direction, an isomorphic version of the Busemann–Petty problem with one measure assigned to each body was considered in [7, Theorem 1], where a comparison between measures was established in terms of the outer volume ratio distance to the class of generalized intersection bodies.
@@ 2. Preliminaries
We follow [3, 6] for standard definitions concerning the Fourier transform of distributions. We denote by $\mathcal{S}$ the Schwartz space of complex-valued test functions on $\mathbb{R}^{n}$—that is, $C^{\infty}$ functions decreasing faster than any polynomial together with all their derivatives.
@@
A distribution $f$ is called positive if $\langle f,\varphi\rangle\geq 0$ for every nonnegative $\varphi\in\mathcal{S}$. It is called positive-definite if $\hat{f}$ is positive. Moreover, any positive-definite distribution is the Fourier transform of a nonnegative tempered measure (see [3, 6]).
@@
The Fourier transform of an even homogeneous distribution of degree $p\in(-n,0)$ is again an even homogeneous distribution of degree np-n-p (see [6]).
Let $f$ be a homogeneous function of degree $n+p\in(-n,0)$ which is also a positive-definite distribution. Then the Fourier transform $\hat{f}$ is a tempered measure that can be expressed in polar coordinates (see [6, Lemma 2.24]). Hence there exists a nonnegative measure $\gamma_{0}$ on $S^{n-1}$ such that
@@
The following proposition, known as the spherical Parseval formula, corresponds to [6, Corollary 3.23] with $k=1$.
@@
The relation between the Fourier and Radon transforms is given by (see [6, Lemma 3.7])